[Brahms-l] recent shifts

From: flemming videbaek <videbaek@rcf.rhic.bnl.gov>
Date: Mon Jan 12 2004 - 16:58:02 EST
Hi
I noticed by inspectrion of the recent supermonito outpuit plots

a) DVM T1, T2 do not show any signal. Is HV on ??
b) as shift report H2 slat 25

Otherwisae it looked ok.

Too the ntuple plots should be revisited e.g since the ZDc are in spec triggers we rarely if at all run
bb triggers, some plots requires those.

Flemming

----------------------------------------------------------------
Flemming Videbaek
Physics Department
Brookhaven National Laboratory

e-mail: videbaek@bnl.gov
phone: 631-344-4106



_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Mon Jan 12 16:49:36 2004

This archive was generated by hypermail 2.1.8 : Mon Jan 12 2004 - 16:49:51 EST